OpenStudy (anonymous):

Find the quotient of the complex numbers. Express your answer in trigonometric form. z1 = 7(cos 15degrees + i sin 15degrees) z2 = 2(cos 110degrees + i sin 110degrees) A. 7/2 (cos 265degrees + i sin 265degrees) B. 5 (cos (-95degrees) + i sin (-95degrees)) C. 14 (cos 125degrees + i sin 125degrees) D. 9 (cos 125degrees + i sin 125degrees)

OpenStudy (anonymous):

Here's the formula: \[\frac{ r _{1} }{ r_{2} } (\cos [\theta _{1} - \theta _{2}] + i \sin [\theta _{1} - \theta _{2}])\]

OpenStudy (anonymous):

When you look at it, it could possibly be A because of 7/2. But if you subtract the degrees, it would look like B. I need help!

OpenStudy (anonymous):

It is A everyone :)